OpenStudy (anonymous):

Emily wants to hang a painting in a gallery but the painting and frame must have an area of 31 square feet. The painting is 5 feet wide by 6 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined?

OpenStudy (anonymous):

4x2 + 22x − 1 = 0 4x2 + 22x + 31 = 0 x2 + 11x − 1 = 0 x2 + 11x + 31 = 0

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

(5 + 2x)(6 + 2x) = 31 expand that

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

x2 + 11x + 31 = 0

OpenStudy (lgbasallote):

how?

OpenStudy (anonymous):

4x2 + 22x + 31 = 0

OpenStudy (lgbasallote):

okay..i think you're guessing

OpenStudy (lgbasallote):

just FOIL..no need to guess

OpenStudy (anonymous):

i believe it's 4x2 + 22x − 1 = 0

OpenStudy (anonymous):

im right ?

OpenStudy (lgbasallote):

did you guess?

OpenStudy (anonymous):

no not this time

OpenStudy (lgbasallote):

well anyways it's right

OpenStudy (anonymous):

thank you :)

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

medal please

OpenStudy (lgbasallote):

you want me to medal you?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

heh just this once. fine.

OpenStudy (anonymous):

thank you